OpenStudy (liizzyliizz):

estimate the Limit for x-2/x^2+x-6

OpenStudy (zarkon):

as x goes to?

OpenStudy (liizzyliizz):

as x goes to 1. lol sorry i just realized i forgot to add that

OpenStudy (zarkon):

you need to make a table?

OpenStudy (zarkon):

you could just put in the number 1 for x to get the exact value

OpenStudy (liizzyliizz):

yeah i know, i just wasnt sure if the values needed have to correspond to the 1? (sorry its my first day learning this and my AP calculus teacher is like insane o.o 0

OpenStudy (liizzyliizz):

the limit should be 4 then?

OpenStudy (zarkon):

no..not is the function is \[\frac{x-2}{x^2+x-6}\]

OpenStudy (zarkon):

or is it \[x-\frac{2}{x^2}+x-6\]

OpenStudy (zarkon):

or something else

OpenStudy (liizzyliizz):

hmm ok so what am i doing wrong.. uhh my teacher explained it weird. i factored out the denominator got (x-2)(X+3) cancelled out the x-2 whic creates a hole in the graph. and used the x+3...

OpenStudy (liizzyliizz):

umm no you were right the first time

OpenStudy (zarkon):

then you would get \[\frac{1}{4}\]

OpenStudy (zarkon):

\[\frac{1-2}{1^2+1-6}=\frac{-1}{-4}=\frac{1}{4}\]

OpenStudy (liizzyliizz):

ok so im going to try the table now umm would i use points like .o1, .001 an stuff like that? thats what my teacher was using in different problems. so i was like o.o

OpenStudy (zarkon):

no

OpenStudy (zarkon):

you are going to 1 not zero

OpenStudy (zarkon):

.9 .99 .999 .9999

OpenStudy (zarkon):

1.1 1.01 1.001 1.0001

OpenStudy (liizzyliizz):

oh ok, i see what you are saying, it just depends on what x is going yo?

OpenStudy (zarkon):

yes

OpenStudy (liizzyliizz):

ahh see you make this much more clearer then my math teacher thank you so much :D

OpenStudy (zarkon):

no problem